OpenStudy (anonymous):

{a^(1/2)+x^(1/2)}/{a^(1/2)-x^(1/2)} Find dy/dx.

OpenStudy (amistre64):

\[(a^{.5}+x^{.5})~(a^{.5}-x^{.5})^{-1}\] use the product rule

OpenStudy (amistre64):

a^.5 is just some constant .... let it be 3 if need be

OpenStudy (amistre64):

\[([3]+x^{.5})~([3]-x^{.5})^{-1}\] \[([3]+x^{.5})'~([3]-x^{.5})^{-1}+([3]+x^{.5})~([3]-x^{.5})'^{-1}\] \[\frac{1}{2x^{.5}}'~([3]-x^{.5})^{-1}+([3]+x^{.5})~([3]-x^{.5})'^{-1}\] \[\frac{1}{2x^{.5}}'~([3]-x^{.5})^{-1}-\frac{1}{2x^{.5}}([3]+x^{.5})~([3]-x^{.5})^{-2}\]

OpenStudy (amistre64):

got a spurious little ' that i didnt edit

OpenStudy (amistre64):

and i forgot the - from the last one .... \[\frac{1}{2x^{.5}}~([3]-x^{.5})^{-1}+\frac{1}{2x^{.5}}([3]+x^{.5})~([3]-x^{.5})^{-2}\]